A uniform solid disk and a uniform hoop are placed
Answer #1

(a) the uniform disk has a lower moment of inertia and arrives first.

(b) with no sliping taking place, the two objects will be in rotation and translation.

energy is conserved

\Delta Ki = -\Delta P

1/2M\upsilon ^2 + 1/2I\omega ^2 = Mgh

M\upsilon ^2 +I*\upsilon ^2/R^2 = 2Mgh

For The Disk I=MR^2/2 and

for the hoop I=MR^2

We finally get \upsilon = \sqrt{4/3gh} for the disk

and \upsilon = \sqrt{gh}   for the hoop.